What is the probability of getting 1000 out of a roll from 1 to 1000 and u have 40 trys.
Answer by stanbon(75887)

What is the probability of getting 1000 out of a roll from 1 to 1000 and u have 40 trys.
------------------
Binomial Problem with n=40, p = 1/1000, x= 1
---
P(x=1) = 40C1(1/1000)^1(999/1000)^39 = 40*(999^39)/(1000)^40 = 0.0385
==========================================================================
Cheers,
Stan H.